# On the Fundamental Impossibility of Hallucination Control in Large Language Models

### v5

@@ 7.5 The Outer Limits of Reason
7.6 Dennett’s Multiple Drafts Model
@@ 7.6.1 Multiple Drafts as Competing Agents
In our transformer analysis in Section 6, each attention head generating logit vector $\mathbf{l}^{(h)}$ writes its own draft of the next token. When Head 1 votes for fox while Head 2 votes for dog, we observe as competing narratives bidding for selection.
7.6.2 Formation of Artificial Consciousness
@@
By Theorem 8 and 9, the Jensen gap quantifies the excess confidence in reports that cannot be justified by constituent drafts. Without that overconfidence a victorious idea that reached the external world (as an LLM response) could loose the fierce competition with other ideas and remain inaccessible or hidden in the knowledge space.
Rigorous and careful neuro-scientific studies of the evolution of brain and intelligence, summarized by Max Bennet in bennett2023brief , show that at some point in history our brains gained the ability to model others’ minds or predict what others may be thinking. That ability provided great advantages in communication and collecting information for future use. Understanding what someone else may think helps formulate context-related questions that result in better answers. When we anticipate understanding of a message we formulate, we have a much better chance to make the knowledge hidden in someone else’s brain accessible for us. Therefore, as suggested by Dennett in Dennett1991-DENCE , it is plausible that sometime in the past someone, while being alone, did accidentally hear his or hers own question fine-tuned to get into another mind, and that triggered generation of an answer in the very same mind asking the question. That answer generation process could then have made some relevant but previously hidden knowledge accessible to the curious mind unaware of its possessions. The hunger for uncertainty reduction could have been satisfied as a result of that inner dialog pushing ideas to bid for attention and conscious reception as a final thought expressed in a language.
@@
7.6.3 A Critique
Dennett’s model has been criticized by many, and it is beyond the scope of this work to present that critique in its full capacity. Instead, let us only refer to the two basic points and speculate how the availability of LLMs could help philosophers refute available theories.
First, Searle john1997mystery and others argue that Dennett eliminates consciousness rather than explaining it, arguing that the spread of ideas is not driven by blind random forces but requires intentionality. The compelling and misleading nature of LLM responses that are generated by random sampling of trained distributions strongly challenge that line of thinking. This paper suggests that the experiments focusing on LLM-hallucination detection may prove useful here to measure or identify intentionality (once we can have its programmable definition).
Second, there has been a charge that the multiple drafts model makes no concrete, falsifiable predictions, rendering it difficult to verify or integrate with neuroscience. This paper provides a mathematical toolkit for experimental philosophy to make that and other models a testable scientific hypothesis in the LLM-based laboratory.
@@ 8 Summary
We introduced a rigorous mathematical framework and grounded in it the general analysis of inference processes and their LLM implementations. At the heart of our mathematical framework lies the semantic information measure $\mu_{C}$ and the emergence operator $\mathcal{E}_{C}$. The earlier one is a context-dependent metric capturing how knowledge reduces uncertainty within computational bounds. The later one formalizes how reasoning makes latent knowledge explicit rather than creating it anew. Finally, we model LLM inference as an auction of ideas–a idealized marketplace where neural components compete by bidding with their partial knowledge, each trying to influence the final response.
@@
These mathematical insights we have made yield intriguing philosophical implications. Hallucination and creativity (two sides of the same coin we call imagination) emerge as mathematically identical phenomena, distinguished only by our normative judgments. The impossibility result explores the outer limits of reason, identified by Gödel’s incompleteness in mathematics, Heisenberg’s uncertainty in physics, and Arrow’s impossibility in social choice. Also, our framework seems to provide mathematical foundations and experimental setting for Dennett’s Multiple Drafts model of consciousness, suggesting that the measurable and observable excess confidence of LLM responses might be one of signatures of conscious processing of a primitive form.
Our goal was also to develop a theoretical framework that may open interesting avenues for future research. Some of ideas and bold speculations competing for attention can be mentioned here. In mathematical consciousness studies, we might formalize the notion of cognitive hunger as an attractor in knowledge space, sustained by continuous violations of information conservation. This could enable us to measure artificial consciousness signatures and test whether biological systems exhibit similar knowledge conservation violations. The framework suggests artificial consciousness might be studied as a nonlinear dynamical system whose attractors emerge from information creation events that our theorems describe. And so, the suggested implications may extend beyond language models to fundamental questions about intelligence itself.

### v7

@@ 8.6 The Outer Limits of Reason
8.7 Formation of Artificial Consciousness
@@
In our transformer analysis in Section 6, each attention head generating logit vector $\mathbf{l}^{(h)}$ writes its own draft of the next token. When Head 1 votes for fox while Head 2 votes for dog, we observe as competing narratives bidding for selection. Then the feedforwad block edits the draft by updating the residual stream.
@@
By Theorem 8 and 9, the Jensen gap quantifies the excess confidence in reports that cannot be justified by constituent drafts. Without that excess confidence a victorious idea cannot win the fierce competition with other ideas. It remains inaccessible or hidden in the knowledge space.
Rigorous and careful neuro-scientific studies of the evolution of brain and intelligence, summarized by Max Bennet in bennett2023brief , show that at some point in history our brains gained the ability to model others’ minds or predict what others may be thinking. That ability provided great advantages in communication and collecting information for future use. Understanding what someone else may think helps formulate context-related questions that result in better answers. When we anticipate understanding of a message we formulate, we have a much better chance to make the knowledge hidden in someone else’s brain accessible for us. Therefore, as suggested by Dennett in Dennett1991-DENCE , it is plausible that sometime in the past someone did accidentally hear his or hers own question fine-tuned to get into another mind, and that triggered generation of an answer in the very same mind asking the question. That answer generation process could then have made some relevant but previously hidden knowledge accessible to the curious mind unaware of its own possessions. The hunger for uncertainty reduction could have been satisfied as a result of that inner dialog pushing ideas to bid for attention and conscious reception as a final thought expressed in a language.
@@
In fact, to close that section, why don’t we situate the framework of Auction of Ideas within the established landscape of consciousness research, as provided by Seth and Bayne in seth2021being . The landscape distinguishes between Global Workspace Theories (GWTs), Predictive Processing (PP) theories, Higher-Order Theories (HOTs), and Integrated information theories (IIT). The Auction of Ideas provides a unifying mechanism that bridges several of these perspectives, but the connection to PP is particularly interesting. The theory characterizes conscious experience as a controlled hallucination, i.e., an optimal inference or best guess about the causes of sensory inputs. Our framework provides a mathematical grounding for this constructive process.
Theorem 8 and Theorem 11 prove that achieving Knowledge-Constrained Optimality, i.e., the best guess, requires a violation of the Semantic Information Conservation. This violation could be interpreted as a mathematical signature of imagination or hallucination, and provides a mathematical analogue to the PP claim: the constructive nature of experience inherent in the best guess is a necessity arising from the aggregation of distributed evidence.
@@ 9 Summary
We introduced a rigorous mathematical framework and grounded in it the general analysis of inference processes and their LLM implementations. At the heart of our mathematical framework lies the semantic information measure $\mu_{C}$ and the emergence operator $\mathcal{E}_{C}$. The earlier one is a context-dependent metric capturing how knowledge reduces uncertainty within computational bounds. The later one formalizes how reasoning makes latent knowledge explicit rather than creating it anew. Finally, we model LLM inference as an auction of ideas—a idealized marketplace where neural components compete by bidding with their partial knowledge, each trying to influence the final response.
@@
These mathematical insights we have made yield some interesting philosophical implications. Hallucination and creativity (two sides of the same coin we call imagination) emerge as mathematically identical phenomena, distinguished only by our normative judgments. The impossibility result explores the outer limits of reason, identified by Gödel’s incompleteness in mathematics, Heisenberg’s uncertainty in physics, and Arrow’s impossibility in social choice. Also, our framework seems to provide mathematical foundations and experimental setting for models of artificial consciousness. That may open interesting avenues for future research.